# Object Segmentation-Assisted Inter Prediction for Versatile Video Coding

### v1

@@ VII-B2 Complexity
| Name | Y | Y | Y | Y | U | V | Y | U | V | Y | U | V | ||||||||||||||
| MarketPlace | - | -18.51% | -2.50% | - | - | - | -13.80% | 0.31% | 0.12% | -4.25% | -3.40% | -3.37% | ||||||||||||||
| RitualDance | - | -25.78% | -2.28% | - | - | - | -7.38% | 21.24% | 21.23% | -2.23% | -2.13% | -2.00% | ||||||||||||||
| Cactus | - | -21.26% | -1.24% | -0.90% | -0.70% | -0.90% | -15.97% | 1.45% | 4.13% | -4.39% | -4.08% | -4.00% | ||||||||||||||
| BasketballDrive | -0.18% | -25.95% | -1.98% | -0.90% | -0.80% | -1.10% | -13.14% | 13.33% | 11.14% | -3.17% | -3.29% | -2.89% | ||||||||||||||
| BQTerrace | - | -23.45% | -0.78% | -1.50% | -1.20% | -1.30% | -23.84% | -8.67% | -10.70% | -2.13% | -1.62% | -1.55% | ||||||||||||||
| Kimono | +0.04% | -19.11% | -0.82% | -1.10% | -1.30% | -0.80% | -7.64% | 14.96% | 13.62% | -2.23% | -2.78% | -2.32% | ||||||||||||||
| ParkScene | - | -16.45% | -0.57% | -1.10% | -1.20% | -0.70% | -11.18% | 7.55% | 9.02% | -1.71% | -1.52% | -1.21% | ||||||||||||||
| BasketballDrill | - | -31.04% | -3.00% | -1.40% | -2.10% | -1.80% | -12.05% | 12.68% | 10.29% | -2.70% | -2.99% | -3.24% | ||||||||||||||
| BQMall | - | -24.41% | -3.69% | -1.60% | -1.40% | -1.60% | -10.11% | 13.54% | 13.47% | -4.38% | -4.87% | -5.01% | ||||||||||||||
| PartyScene | - | -20.51% | -1.50% | -1.50% | -1.00% | -0.70% | -5.46% | 17.11% | 17.49% | -1.68% | -1.94% | -1.93% | ||||||||||||||
| RaceHorsesC | -1.05% | -14.49% | -2.60% | -0.70% | -0.50% | -0.40% | -10.85% | 18.34% | 19.93% | -2.58% | -4.09% | -4.42% | ||||||||||||||
| BasketballPass | - | -17.59% | -2.79% | -0.90% | -1.10% | -1.00% | -8.09% | 18.51% | 17.10% | -1.66% | -2.28% | -2.27% | ||||||||||||||
| BQSquare | - | -23.14% | -2.64% | -1.60% | -2.40% | -2.20% | -4.37% | 5.95% | 7.04% | -2.17% | -2.07% | -1.50% | ||||||||||||||
| BlowingBubbles | -0.17% | -14.15% | -2.27% | -1.80% | -1.90% | -2.10% | -7.52% | 15.82% | 17.49% | -2.43% | -2.00% | -2.66% | ||||||||||||||
| RaceHorses | - | -14.06% | -2.99% | -0.80% | -0.70% | -0.70% | -6.52% | 20.78% | 18.72% | -2.05% | -3.04% | -2.75% | ||||||||||||||
| FourPeople | - | -26.19% | -1.26% | -0.80% | -0.60% | -0.90% | -14.32% | -6.64% | -8.85% | -1.82% | -1.37% | -1.29% | ||||||||||||||
| Johnny | - | -28.65% | -3.38% | -0.80% | -0.50% | -0.20% | -16.01% | -5.92% | -10.13% | -3.41% | -2.80% | -3.03% | ||||||||||||||
| KristenAndSara | - | -24.97% | -1.62% | -1.00% | -0.70% | -0.90% | -14.19% | -4.63% | -5.28% | -2.07% | -1.63% | -1.56% | ||||||||||||||
| Tango2 | - | - | - | -0.80% | -1.10% | -1.20% | -11.59% | 4.54% | 3.47% | -2.90% | -3.32% | -2.85% | ||||||||||||||
| Drums100 | - | - | - | -1.40% | -1.80% | -1.60% | -17.93% | -2.46% | -1.84% | -1.78% | -1.93% | -1.23% | ||||||||||||||
| Campfire | - | - | - | -1.70% | -2.50% | -3.30% | -20.34% | 9.60% | 8.37% | -0.72% | -0.47% | -0.88% | ||||||||||||||
| ToddlerFountain2 | - | - | - | -0.80% | -0.90% | -1.30% | -6.43% | 48.16% | 38.69% | -0.23% | -1.59% | -0.57% | ||||||||||||||
| TrafficFlow | - | - | - | -0.90% | -1.10% | -0.90% | -14.30% | 2.16% | 2.74% | -1.44% | -1.35% | -1.29% | ||||||||||||||
| DaylightRoad2 | - | - | - | -1.10% | -1.00% | -1.10% | -17.23% | -8.60% | -7.49% | -3.36% | -2.47% | -2.54% | ||||||||||||||
| FoodMarket4 | - | - | - | - | - | - | -9.80% | 10.05% | 11.89% | -1.83% | -1.81% | -1.21% | ||||||||||||||
| CatRobot | - | - | - | - | - | - | -15.13% | 0.94% | 1.21% | -4.24% | -5.45% | -5.28% | ||||||||||||||
| ParkRunning3 | - | - | - | - | - | - | -17.23% | 9.61% | 12.50% | -3.92% | -4.07% | -4.32% | ||||||||||||||
| Average | -0.34% | -22.14% | -2.10% | -1.14% | -1.20% | -1.22% | -12.31% | 8.51% | 7.98% | -2.50% | -2.61% | -2.49% | ||||||||||||||
Note: The indication of table notes , , , is the same as the table notes of Table III .
VII-B3 Comparisons with the Other State-of-the-Art Methods
@@
Table III and IV show the experimental data comparing our proposed SAIP with the previous related work on segmentation-based partition methods [13] and [14], under LDP, LDB and RA configurations ([13] did not reported RA configuration, [14] did not reported LDP configuration). These related methods use different anchors. Specifically, [13] and [14] use HM-14.0 and VTM-1.0 as the anchor, respectively. We directly cite the results reported in [13] and [14], since our reimplementation may not work as well as the authors’. HM-14.0[44] and VTM-1.0[45] do not contain some new technical tools in the latest version of VVC, which have a large performance gap with VTM-12.0. To make a fair comparison with related work, we test SAIP in a simplified VTM-12.0 with some coding tools turned off, and compare it with [13] [14] on HEVC and VVC test sequences. Note that the turned-off coding tools in simplified VTM-12.0 are all not supported in HM-14.0[44] and VTM-1.0[45]. (The number of coded frames of these methods in Table III and IV follows the number of coded frames specified in the Common Test Condition of HEVC[46] and VVC[42].) From Table III and IV , it is observed that the proposed SAIP achieves a better relative BD-rate reduction on higher version than the other methods on low versions, which demonstrates the superior performance of the proposed method.
@@
| Low-delay P Configuration (%) | ||||||||||||
| Class | ONE-STEP SA-MC | TWO-STEP SA-MC | w/o SA-MVC | SAIP (SA-MC+SA-MVC) | ||||||||
| Y | U | V | Y | U | V | Y | U | V | Y | U | V | |
| B | -0.05% | 0.03% | -0.29% | -0.10% | -0.17% | -0.11% | -0.11% | -0.48% | -0.20% | -0.25% | -0.21% | -0.45% |
| C | -0.52% | -0.64% | -0.69% | -0.89% | -1.21% | -0.70% | -1.07% | -1.23% | -1.31% | -1.26% | -1.43% | -1.41% |
| E | -0.49% | -0.51% | -0.31% | -0.68% | -0.90% | -0.74% | -0.64% | -1.46% | -0.48% | -1.17% | -1.19% | -0.48% |
| Avg. | -0.32% | -0.33% | -0.43% | -0.51% | -0.70% | -0.46% | -0.56% | -0.98% | -0.64% | -0.82% | -0.86% | -0.78% |
| D | -0.58% | -0.52% | -0.93% | -0.83% | -0.89% | -1.28% | -0.60% | -0.64% | -0.64% | -0.81% | -1.01% | -0.10% |
| Low-delay B Configuration (%) | ||||||||||||
| Class | ONE-STEP SA-MC | TWO-STEP SA-MC | w/o SA-MVC | SAIP (SA-MC+SA-MVC) | ||||||||
| Y | U | V | Y | U | V | Y | U | V | Y | U | V | |
| B | 0.01% | -0.08% | 0.14% | -0.11% | -0.13% | 0.00% | -0.12% | -0.37% | 0.11% | -0.30% | -0.08% | -0.17% |
| C | -0.07% | 0.00% | -0.29% | -0.42% | -0.79% | -0.08% | -0.32% | -0.37% | -0.75% | -0.60% | -0.58% | -0.37% |
| E | -0.33% | -0.12% | -0.37% | -0.43% | -0.62% | -0.20% | -0.27% | -0.45% | -0.07% | -0.66% | -0.84% | 0.25% |
| Avg. | -0.10% | -0.06% | -0.13% | -0.29% | -0.47% | -0.08% | -0.22% | -0.39% | -0.22% | -0.49% | -0.45% | -0.13% |
| D | -0.37% | -0.65% | 0.45% | -0.47% | 0.14% | 0.35% | -0.12% | 0.21% | 0.93% | -0.41% | -0.49% | 0.39% |
| Random Access Configuration (%) | ||||||||||||
| Class | ONE-STEP SA-MC | TWO-STEP SA-MC | w/o SA-MVC | SAIP (SA-MC+SA-MVC) | ||||||||
| Y | U | V | Y | U | V | Y | U | V | Y | U | V | |
| A | 0.80% | 0.31% | 0.41% | -0.06% | -0.13% | -0.13% | -0.08% | 0.02% | 0.02% | -0.26% | -0.07% | -0.19% |
| B | 0.76% | 0.55% | 0.65% | -0.05% | 0.07% | -0.10% | -0.09% | -0.04% | -0.04% | -0.26% | 0.21% | -0.24% |
| C | 0.55% | 0.34% | 0.63% | -0.20% | -0.24% | -0.26% | -0.18% | -0.30% | -0.03% | -0.50% | -0.46% | -0.29% |
| E | 0.40% | 0.26% | 0.29% | -0.06% | 0.16% | 0.09% | -0.13% | -0.10% | 0.07% | -0.52% | -0.46% | -0.24% |
| Avg. | 0.67% | 0.38% | 0.51% | -0.09% | -0.05% | -0.11% | -0.11% | -0.09% | 0.00% | -0.37% | -0.14% | -0.23% |
| D | 0.70% | 0.67% | 0.40% | -0.12% | 0.52% | 0.12% | -0.09% | 0.27% | -0.06% | -0.45% | 0.02% | -0.29% |
VII-B4 Overall Performance on the Specific Scenes
@@
According to the performance of SAIP under CTC and the comparison with other methods, the proposed object-aware prediction framework is beneficial to the scene with moving objects and complex motion fields (such as the walkers of BQMall shown in Fig . 7 and the athletes of BasketballDrill), and achieves high performance on these scenes. To further verify the characteristics and advantages of SAIP, we test SAIP on some selected test sequences of specific scenes (diverse scenes, multiple moving objects, complex motions) from previous standards’ CTC [1, 7] and popular datasets [47]. Table VI shows the performance of partial sequences on these specific scenes, the configuration and whole results are detailed in Section IV of supplementary material. We can see that our proposed SAIP can achieve up to 2.90%, 1.37%, 1.02%, and on average 1.01%, 0.78%, 0.52% BD-rate reduction (Y component) on VTM-12.0. The results demonstrate the SAIP can achieve accurate object-aware prediction on specific scenes with more moving objects, and further improve the flexibility of motion modeling in VVC.
| Sequence | Resolution | LDP (%) | LDB (%) | RA (%) |
| CarPhone | QCIF | -1.67% | -1.05% | -0.80% |
| Slient | CIF | -2.90% | -1.37% | -1.02% |
| Vidyo3 | 720P | -1.27% | -0.80% | -0.63% |
| Sunflower | 1080P | -0.60% | -0.64% | -0.49% |
| Narrator | 4K | -1.31% | -1.08% | -0.34% |
| Average (20) | - | -1.01% | -0.78% | -0.52% |
VII-C Performance Analysis

### v2

@@ IX-B2 Complexity
| Name | Y | Y | Y | Y | U | V | Y | U | V | Y | U | V | |||||||||||||||||||
| MarketPlace | - | -18.51% | -2.50% | - | - | - | -13.80% | 0.31% | 0.12% | -4.25% | -3.40% | -3.37% | |||||||||||||||||||
| RitualDance | - | -25.78% | -2.28% | - | - | - | -7.38% | 21.24% | 21.23% | -2.23% | -2.13% | -2.00% | |||||||||||||||||||
| Cactus | - | -21.26% | -1.24% | -0.90% | -0.70% | -0.90% | -15.97% | 1.45% | 4.13% | -4.39% | -4.08% | -4.00% | |||||||||||||||||||
| BasketballDrive | -0.18% | -25.95% | -1.98% | -0.90% | -0.80% | -1.10% | -13.14% | 13.33% | 11.14% | -3.17% | -3.29% | -2.89% | |||||||||||||||||||
| BQTerrace | - | -23.45% | -0.78% | -1.50% | -1.20% | -1.30% | -23.84% | -8.67% | -10.70% | -2.13% | -1.62% | -1.55% | |||||||||||||||||||
| Kimono | +0.04% | -19.11% | -0.82% | -1.10% | -1.30% | -0.80% | -7.64% | 14.96% | 13.62% | -2.23% | -2.78% | -2.32% | |||||||||||||||||||
| ParkScene | - | -16.45% | -0.57% | -1.10% | -1.20% | -0.70% | -11.18% | 7.55% | 9.02% | -1.71% | -1.52% | -1.21% | |||||||||||||||||||
| BasketballDrill | - | -31.04% | -3.00% | -1.40% | -2.10% | -1.80% | -12.05% | 12.68% | 10.29% | -2.70% | -2.99% | -3.24% | |||||||||||||||||||
| BQMall | - | -24.41% | -3.69% | -1.60% | -1.40% | -1.60% | -10.11% | 13.54% | 13.47% | -4.38% | -4.87% | -5.01% | |||||||||||||||||||
| PartyScene | - | -20.51% | -1.50% | -1.50% | -1.00% | -0.70% | -5.46% | 17.11% | 17.49% | -1.68% | -1.94% | -1.93% | |||||||||||||||||||
| RaceHorsesC | -1.05% | -14.49% | -2.60% | -0.70% | -0.50% | -0.40% | -10.85% | 18.34% | 19.93% | -2.58% | -4.09% | -4.42% | |||||||||||||||||||
| BasketballPass | - | -17.59% | -2.79% | -0.90% | -1.10% | -1.00% | -8.09% | 18.51% | 17.10% | -1.66% | -2.28% | -2.27% | |||||||||||||||||||
| BQSquare | - | -23.14% | -2.64% | -1.60% | -2.40% | -2.20% | -4.37% | 5.95% | 7.04% | -2.17% | -2.07% | -1.50% | |||||||||||||||||||
| BlowingBubbles | -0.17% | -14.15% | -2.27% | -1.80% | -1.90% | -2.10% | -7.52% | 15.82% | 17.49% | -2.43% | -2.00% | -2.66% | |||||||||||||||||||
| RaceHorses | - | -14.06% | -2.99% | -0.80% | -0.70% | -0.70% | -6.52% | 20.78% | 18.72% | -2.05% | -3.04% | -2.75% | |||||||||||||||||||
| FourPeople | - | -26.19% | -1.26% | -0.80% | -0.60% | -0.90% | -14.32% | -6.64% | -8.85% | -1.82% | -1.37% | -1.29% | |||||||||||||||||||
| Johnny | - | -28.65% | -3.38% | -0.80% | -0.50% | -0.20% | -16.01% | -5.92% | -10.13% | -3.41% | -2.80% | -3.03% | |||||||||||||||||||
| KristenAndSara | - | -24.97% | -1.62% | -1.00% | -0.70% | -0.90% | -14.19% | -4.63% | -5.28% | -2.07% | -1.63% | -1.56% | |||||||||||||||||||
| Tango2 | - | - | - | -0.80% | -1.10% | -1.20% | -11.59% | 4.54% | 3.47% | -2.90% | -3.32% | -2.85% | |||||||||||||||||||
| Drums100 | - | - | - | -1.40% | -1.80% | -1.60% | -17.93% | -2.46% | -1.84% | -1.78% | -1.93% | -1.23% | |||||||||||||||||||
| Campfire | - | - | - | -1.70% | -2.50% | -3.30% | -20.34% | 9.60% | 8.37% | -0.72% | -0.47% | -0.88% | |||||||||||||||||||
| ToddlerFountain2 | - | - | - | -0.80% | -0.90% | -1.30% | -6.43% | 48.16% | 38.69% | -0.23% | -1.59% | -0.57% | |||||||||||||||||||
| TrafficFlow | - | - | - | -0.90% | -1.10% | -0.90% | -14.30% | 2.16% | 2.74% | -1.44% | -1.35% | -1.29% | |||||||||||||||||||
| DaylightRoad2 | - | - | - | -1.10% | -1.00% | -1.10% | -17.23% | -8.60% | -7.49% | -3.36% | -2.47% | -2.54% | |||||||||||||||||||
| FoodMarket4 | - | - | - | - | - | - | -9.80% | 10.05% | 11.89% | -1.83% | -1.81% | -1.21% | |||||||||||||||||||
| CatRobot | - | - | - | - | - | - | -15.13% | 0.94% | 1.21% | -4.24% | -5.45% | -5.28% | |||||||||||||||||||
| ParkRunning3 | - | - | - | - | - | - | -17.23% | 9.61% | 12.50% | -3.92% | -4.07% | -4.32% | |||||||||||||||||||
| Average | -0.34% | -22.14% | -2.10% | -1.14% | -1.20% | -1.22% | -12.31% | 8.51% | 7.98% | -2.50% | -2.61% | -2.49% | |||||||||||||||||||
Note: The indication of table notes 1, 2, 3, 4 is the same as in Table V .
| Sequence | Resolution | Low-Delay P (%) | Low-Delay B (%) | Random Access (%) | ||||||||||
| Y | U | V | Y | U | V | Y | U | V | ||||||
| Carphone | QCIF | -1.67% | -0.09% | 2.57% | -1.05% | -1.85% | 0.63% | -0.80% | -1.33% | -0.26% | ||||
| Grandma | QCIF | -1.26% | -1.91% | -0.72% | -1.27% | -0.90% | -1.41% | -0.18% | -0.56% | -0.08% | ||||
| Deadline | CIF | -0.92% | -0.81% | -1.89% | -0.81% | -0.44% | 0.81% | -0.67% | -0.63% | -1.19% | ||||
| Akiyo | CIF | -0.69% | 0.05% | -0.50% | -0.86% | -1.46% | -1.02% | -0.01% | 0.24% | 0.41% | ||||
| Silent | CIF | -2.90% | -4.03% | -2.78% | -1.37% | -1.26% | -0.62% | -1.02% | -0.45% | -0.57% | ||||
| Bowing | CIF | -1.71% | -3.15% | -1.58% | -1.08% | -0.48% | -1.40% | -0.72% | -0.37% | -0.69% | ||||
| Football | CIF | -0.71% | 1.39% | -2.06% | -0.70% | -1.24% | 0.38% | -0.63% | 0.83% | -0.47% | ||||
| Foreman | CIF | -0.85% | 0.60% | -2.98% | -0.52% | -1.96% | 1.11% | -0.38% | -0.38% | -0.80% | ||||
| Vidyo1 | 720P | -0.75% | -0.19% | -1.06% | -0.81% | 0.16% | -0.39% | -0.58% | -0.75% | -0.51% | ||||
| Vidyo3 | 720P | -1.27% | -1.80% | -0.88% | -0.80% | -1.22% | -0.29% | -0.63% | -0.45% | -0.33% | ||||
| Vidyo4 | 720P | -0.84% | -0.07% | 0.65% | -0.55% | -0.59% | -0.99% | -0.43% | -0.37% | -0.19% | ||||
| Crew | 720P | -0.73% | 1.25% | -1.20% | -0.59% | -1.48% | -0.98% | -0.52% | -1.82% | 0.51% | ||||
| Aspen | 1080P | -0.51% | -0.20% | -0.37% | -0.42% | -0.53% | 0.12% | -0.36% | -0.17% | 0.08% | ||||
| Red_kayak | 1080P | -0.50% | -0.30% | 0.86% | -0.51% | -1.86% | 0.61% | -0.47% | 0.20% | -1.02% | ||||
| Sunflower | 1080P | -0.60% | -0.92% | -0.86% | -0.64% | -0.72% | -0.42% | -0.49% | 0.01% | -0.25% | ||||
| FoodMarket2 | 4K | -0.45% | -0.63% | -0.21% | -0.47% | -1.13% | 0.36% | -0.51% | -0.32% | 0.24% | ||||
| BoxingPractice | 4K | -1.22% | -1.72% | -0.70% | -0.90% | -1.67% | -1.15% | -0.78% | -1.23% | -0.55% | ||||
| Crosswalk | 4K | -0.67% | 0.18% | -0.85% | -0.63% | -1.86% | -0.35% | -0.31% | -0.70% | -0.71% | ||||
| Narrator | 4K | -1.31% | -1.68% | -2.15% | -1.08% | -1.66% | -1.02% | -0.34% | 0.18% | 0.10% | ||||
| SquareAndTimelapse | 4K | -0.65% | -0.03% | -0.96% | -0.55% | 0.07% | -0.54% | -0.50% | -0.68% | -0.47% | ||||
| Overall | - | -1.01% | -0.70% | -0.88% | -0.78% | -1.10% | -0.33% | -0.52% | -0.44% | -0.34% | ||||
IX-B3 Comparisons with State-of-the-Art Methods
@@
Table V and VI show the experimental data comparing our proposed SAIP with the previous related work on segmentation-based partition methods [13] and [14], under LDP, LDB and RA configurations ([13] did not report RA configuration, [14] did not report LDP configuration). These related methods use different anchors. Specifically, [13] and [14] use HM-14.0 and VTM-1.0 as the anchor, respectively. We directly cite the results reported in [13] and [14], since our reimplementation may not work as well as the authors’. HM-14.0[48] and VTM-1.0[49] do not contain some new technical tools in the latest version of VVC, which have a large performance gap with VTM-12.0. To make a fair comparison with related work, we test SAIP in a simplified VTM-12.0 with some coding tools turned off, and compare it with [13] [14] on HEVC and VVC test sequences. Note that the turned-off coding tools in simplified VTM-12.0 are all not supported in HM-14.0[48] and VTM-1.0[49]. (The number of coded frames of these methods in Table V and VI follows the number of coded frames specified in the Common Test Condition of HEVC[50] and VVC[46].) From Table V and VI , it is observed that the proposed SAIP achieves a better relative BD-rate reduction on higher version than the other methods on low versions, which demonstrates the superior performance of the proposed method.
@@
In addition, we also compare our proposed method with some enhanced geometric-based partitioning methods [20], under LDB and RA configurations. The comparison result is shown in Table IV . Specifically, [20] uses VTM-8.0 as the anchor, and we directly cite the results reported in [20] (did not report LDP configuration). From Table IV , it is observed that the proposed SAIP achieves a good relative BD-rate reduction on the VTM higher version than the other enhanced geometric-based partitioning methods on the VTM low version, which further verifies the efficiency of the proposed method.
IX-B4 On Specific Scenes
@@
According to the performance of SAIP under CTC and the comparison with other methods, the proposed object-aware prediction framework is beneficial to the scene with moving objects and complex motion fields (such as the walkers of BQMall shown in Fig . 9 and the athletes of BasketballDrill), and achieves high performance on these scenes. To further verify the characteristics and advantages of SAIP, we test SAIP on some selected test sequences of specific scenes (diverse scenes, multiple moving objects, complex motions) from previous standards’ CTC [1, 7] and popular datasets [51]. Table VII shows the performance of partial sequences on these specific scenes. We can see that our proposed SAIP can achieve up to 2.90%, 1.37%, 1.02%, and on average 1.01%, 0.78%, 0.52% BD-rate reduction (Y component) on VTM-12.0. The results demonstrate the SAIP can achieve accurate object-aware prediction on specific scenes with more moving objects, and further improve the flexibility of motion modeling in VVC.
| Low-delay P Configuration (%) | ||||||||||||||||
| Class | ONE-STEP SA-MC | TWO-STEP SA-MC | w/o SA-MVC | SAIP (SA-MC+SA-MVC) | ||||||||||||
| Y | U | V | Y | U | V | Y | U | V | Y | U | V | |||||
| B | -0.05% | 0.03% | -0.29% | -0.10% | -0.17% | -0.11% | -0.11% | -0.48% | -0.20% | -0.25% | -0.21% | -0.45% | ||||
| C | -0.52% | -0.64% | -0.69% | -0.89% | -1.21% | -0.70% | -1.07% | -1.23% | -1.31% | -1.26% | -1.43% | -1.41% | ||||
| E | -0.49% | -0.51% | -0.31% | -0.68% | -0.90% | -0.74% | -0.64% | -1.46% | -0.48% | -1.17% | -1.19% | -0.48% | ||||
| Avg. | -0.32% | -0.33% | -0.43% | -0.51% | -0.70% | -0.46% | -0.56% | -0.98% | -0.64% | -0.82% | -0.86% | -0.78% | ||||
| D | -0.58% | -0.52% | -0.93% | -0.83% | -0.89% | -1.28% | -0.60% | -0.64% | -0.64% | -0.81% | -1.01% | -0.10% | ||||
| Low-delay B Configuration (%) | ||||||||||||||||
| Class | ONE-STEP SA-MC | TWO-STEP SA-MC | w/o SA-MVC | SAIP (SA-MC+SA-MVC) | ||||||||||||
| Y | U | V | Y | U | V | Y | U | V | Y | U | V | |||||
| B | 0.01% | -0.08% | 0.14% | -0.11% | -0.13% | 0.00% | -0.12% | -0.37% | 0.11% | -0.30% | -0.08% | -0.17% | ||||
| C | -0.07% | 0.00% | -0.29% | -0.42% | -0.79% | -0.08% | -0.32% | -0.37% | -0.75% | -0.60% | -0.58% | -0.37% | ||||
| E | -0.33% | -0.12% | -0.37% | -0.43% | -0.62% | -0.20% | -0.27% | -0.45% | -0.07% | -0.66% | -0.84% | 0.25% | ||||
| Avg. | -0.10% | -0.06% | -0.13% | -0.29% | -0.47% | -0.08% | -0.22% | -0.39% | -0.22% | -0.49% | -0.45% | -0.13% | ||||
| D | -0.37% | -0.65% | 0.45% | -0.47% | 0.14% | 0.35% | -0.12% | 0.21% | 0.93% | -0.41% | -0.49% | 0.39% | ||||
| Random Access Configuration (%) | ||||||||||||||||
| Class | ONE-STEP SA-MC | TWO-STEP SA-MC | w/o SA-MVC | SAIP (SA-MC+SA-MVC) | ||||||||||||
| Y | U | V | Y | U | V | Y | U | V | Y | U | V | |||||
| A | 0.80% | 0.31% | 0.41% | -0.06% | -0.13% | -0.13% | -0.08% | 0.02% | 0.02% | -0.26% | -0.07% | -0.19% | ||||
| B | 0.76% | 0.55% | 0.65% | -0.05% | 0.07% | -0.10% | -0.09% | -0.04% | -0.04% | -0.26% | 0.21% | -0.24% | ||||
| C | 0.55% | 0.34% | 0.63% | -0.20% | -0.24% | -0.26% | -0.18% | -0.30% | -0.03% | -0.50% | -0.46% | -0.29% | ||||
| E | 0.40% | 0.26% | 0.29% | -0.06% | 0.16% | 0.09% | -0.13% | -0.10% | 0.07% | -0.52% | -0.46% | -0.24% | ||||
| Avg. | 0.67% | 0.38% | 0.51% | -0.09% | -0.05% | -0.11% | -0.11% | -0.09% | 0.00% | -0.37% | -0.14% | -0.23% | ||||
| D | 0.70% | 0.67% | 0.40% | -0.12% | 0.52% | 0.12% | -0.09% | 0.27% | -0.06% | -0.45% | 0.02% | -0.29% | ||||
IX-C Performance Analysis